Hank, who is twice the mass of Marcia, is running east at 2.0 m/s. Marcia is running north at 1.8 m/s. They collide and stick. Immediately after the collision, what is their speed?
Answer #1

Solution :-

(Vx) = velocity of the x-component

(Vy) = velocity of the y- component

conserving the momentum in x-direction we have

m×2 = (m + 2m)Vx

Vx = 0.667 m/s

conserving the momentum in y-direction we hav e

2m×1.8 = (m + 2m)Vy

Vy = 1.2 m/s

speed = sqrt(0.667^ + 1.2^2) = 1.372m/s

direction = tan^-1 ( 1.2 / 0.667) = 60.93°